AB is a chord of a circle O.If |AB|=24.2cm and the perimeter of triangle AOB is 52.2 cm,calculate angle AOB correct to the nearest degree


1
Expert's answer
2021-06-01T06:51:09-0400

Perimeter of triangle = "52.2\space cm"

Let radius of circle = "R"

"R+R+AB=52.2"

"2R=52.2-AB=28"

"R=14\space cm"

"\angle AOB=2\sin^{-1}\bigg(\dfrac{12.1}{14}\bigg)"

"=2\times59.801\\=119.6\degree"